OpenStudy (chycora):

Jered's parents are going to repaint their house. They need to select two colors to complete the job. If they previously narrowed their decision down to six colors, how many combinations of two can they choose from? A. 12 B. 15 C. 30 D. 3

OpenStudy (chycora):

a or d

OpenStudy (studygurl14):

Neither.

OpenStudy (chycora):

i'm really thinking a

OpenStudy (chycora):

then it's c

OpenStudy (studygurl14):

No, not that either.

OpenStudy (chycora):

ok then i'm confused

OpenStudy (studygurl14):

nCr, that's what you're doing. n = total to choose from, r= number of items you take from the n \(\Large\frac{n!}{(n-r)!r!}\)

OpenStudy (chycora):

how

OpenStudy (studygurl14):

n = 6, r=2, plug in and solve

OpenStudy (kendricklamar2014):

@Agl202 if its not A, C, or D. Well of course its B

OpenStudy (chycora):

i gt 3 so like i said d

OpenStudy (studygurl14):

How did you get 3? @chycora Remember that 6! = 6 x 5 x 4 x 3 x 2 x 1

OpenStudy (chycora):

so when i do that i get 720 that is nt even close to 15 so still don't get it

OpenStudy (studygurl14):

That's not the whole formula. Look u p there^ to see all of it.

OpenStudy (studygurl14):

\(\Large\frac{6!}{(6-2)!2!}\)

OpenStudy (chycora):

ok i did that wel watever i puting in b so thanx

OpenStudy (studygurl14):

you're welcome

OpenStudy (studygurl14):

\(\Large \frac{6!}{(6-2)!2!}\rightarrow \frac{720}{4!(2)}\rightarrow \frac{360}{24}\rightarrow \large 15\)